# Consistent Opponent Modeling in Imperfect-Information Games

### v5

@@ 3 Algorithm
Proposition 4 (Consistency under standard assumptions).
Assume (i) the opponent’s true strategy $\sigma^{*}_{-i}$ lies in the model class and has positive prior probability, (ii) the data-generating process is correctly specified and identifies $\sigma^{*}_{-i}$ (i.e., distinct opponent strategies induce distinct distributions over the observations generated by the algorithm), and (iii) all opponent information sets/actions whose probabilities are being estimated are visited infinitely often (equivalently, the induced observation process provides persistent excitation). Then the algorithm is consistent: the posterior concentrates at $\sigma^{*}_{-i}$ and $\lim_{t\to\infty}M_{t}=\sigma^{*}_{-i}$.
@@
At each iteration $t$ the algorithm computes the modeled strategy $M_{t}$ that maximizes the posterior (MAP). Under assumptions (i)–(iii), standard Bayesian consistency results imply that the posterior distribution concentrates to a point mass at the true opponent strategy $\sigma^{*}_{-i}$ as the number of observations (iterations) goes to infinity. Therefore the posterior mode converges to $\sigma^{*}_{-i}$, i.e., $\lim_{t\rightarrow\infty}M_{t}=\sigma^{*}_{-i}$. ∎
Remark. If some opponent information sets are reached with zero probability under the (possibly exploitative) play induced by the algorithm, then the corresponding parameters are not identifiable from on-policy data and the conclusion can fail.
@@ 4 Experiments
Figure 2: Profit as a function of game iteration for several opponent modeling algorithms and benchmark strategies. The results are averaged over 100 opponents generated randomly from the prior distribution. The sampling algorithms all use 10 samples.

### v6

@@ 3 Algorithm
Proposition 4 (Consistency under Persistent Excitation).
Let $\mathcal{M}$ denote the set of feasible opponent strategies in sequence form. Assume: (i) the true opponent strategy $\boldsymbol{\sigma}^{*}_{-i}$ lies in the interior of $\mathcal{M}$ and the prior assigns it positive density; (ii) the induced observation model is identifiable, i.e., distinct strategies in $\mathcal{M}$ induce distinct distributions over observed play; and (iii) (persistent excitation) every opponent information set is visited infinitely often with nonzero frequency.
Then the posterior distribution concentrates at $\boldsymbol{\sigma}^{*}_{-i}$, and any sequence of MAP estimates $M_{t}$ converges almost surely to $\boldsymbol{\sigma}^{*}_{-i}$.
@@
Let $\ell_{t}(\boldsymbol{\sigma})$ denote the log-likelihood of the observation at iteration $t$ under opponent strategy $\boldsymbol{\sigma}$. By Bayes’ rule,
$\log p(\boldsymbol{\sigma}\mid\mathcal{D}_{t})=\log p(\boldsymbol{\sigma})+\sum_{s=1}^{t}\ell_{s}(\boldsymbol{\sigma}).$
Under assumption (iii), each component of the opponent’s strategy is observed infinitely often, ensuring convergence of empirical log-likelihood averages to their expectations. By identifiability (assumption (ii)), the expected log-likelihood is uniquely maximized at the true strategy $\boldsymbol{\sigma}^{*}_{-i}$. Hence, for any $\boldsymbol{\sigma}\neq\boldsymbol{\sigma}^{*}_{-i}$, the cumulative log-likelihood difference $\sum_{s=1}^{t}[\ell_{s}(\boldsymbol{\sigma})-\ell_{s}(\boldsymbol{\sigma}^{*}_{-i})]$ diverges to -$\infty$.
Since the prior assigns positive density in a neighborhood of $\boldsymbol{\sigma}^{*}_{-i}$, the posterior mass outside any neighborhood of $\boldsymbol{\sigma}^{*}_{-i}$ converges to zero. Therefore, the posterior concentrates at $\boldsymbol{\sigma}^{*}_{-i}$, and any sequence of MAP estimators converges to $\boldsymbol{\sigma}^{*}_{-i}$. ∎
The assumptions underlying Proposition 4 are standard in the literature on Bayesian consistency and learning under partial observability. In particular, the requirement that the true opponent strategy lies in the support of the prior, together with identifiability of the induced observation process, corresponds to classical sufficient conditions for posterior consistency [15, 2, 7]. Identifiability ensures that distinct opponent strategies induce distinguishable distributions over observed play, so that the true strategy can in principle be recovered from infinite data. The persistent excitation assumption plays a role analogous to excitation conditions in adaptive control, ensuring that informative observations occur infinitely often and preventing convergence to uninformative regions of the strategy space [9, 3]. Together, these conditions ensure posterior concentration around the true opponent strategy even when observations are generated endogenously.
@@ 4 Experiments
Figure 2: Expected payoff as a function of game iteration for several opponent modeling algorithms and benchmark strategies. The results are averaged over 100 opponents generated randomly from the prior distribution. The sampling algorithms all use 10 samples.